Grade 12th passMagical Mathematics[Interesting Approach]

Interesting facts on mathematics

jagan
11 Years agoGrade 12th pass

1 Answer

Raheema Javed
11 Years ago
The Sweepstakes Law
Sections 197 to 206 of theCriminal Code of Canada, when mingled provide a ban on “for-profit gaming or betting”, with exceptions like licensedcasinos, provinciallotteries and charity events.However, organizations take advantage of the fact that the law allows prizes to be given forgames requiring skills, or games needing both skill and chance. In order to make such contests official, these games usually consist of a mathematical STQ (Skill Testing Equation). Because of this law, the lottery winners can’t be chosen by luck. If you happen to win the lottery, you will have to answer a skill question, which is usually a four-part mathematical test such as “314+1÷ 5-9

The Cursed 528thDigit
Mathematician William Shanks calculated the value of Pi (π) to 707 places but made a mistake on the 528thdigit and thereby calculating every digit after it incorrectly. Pi (π) is not capable of being expressed as a fraction, thus, making it an irrational number. It neither repeats and nor does it end when written as a decimal.

The Dyscalculia Disease
Now the low grades in the mathematics paper can be explained to the furious parents by the knowledge of a simple disease and good acting skills. The word dyscalculia means difficulty in understanding numbers, and learning mathematics facts. Convincing, isn’t it?

The Four Lucky Ones

There are just four numbers (after 1) which are the sums of the cubes of their digits:

153 = 1^3 + 5^3 + 3^3

370 = 3^3 + 7^3 + 0^3

371 = 3^3 + 7^3 + 1^3

407 = 4^3 + 0^3 + 7^3 [ ^ means raised to power]

The Expensive Zero
In 1997, a divide by zero error occurred when a crew member entered “a” in the database on board of the USS Yorktown, a billion dollar navy smart boat, brought down all the machines on the network causing the ship’s thrust system to fail.